OpenStudy (anonymous):

X^2+6x=7 Can someone explain this? And find a,b, and c?

OpenStudy (anonymous):

x^2+6x-7=0 x^2-x+7x-7=0 x(x-1)+7(x-1)=0 (x+7)(x-1)=0 x=-7,x=1

OpenStudy (anonymous):

I'm just having a bit of a hard time understanding the third step, is there any sort of explanation?

OpenStudy (anonymous):

i just take x common

OpenStudy (anonymous):

one is x square and another is x so you can takke x common from each other

OpenStudy (amistre64):

himalya thinks you are trying to solve for x, as opposed to determining the specific coefficients of the standard: ax^2 + bx + c